# Mutual Information Density of Massive MIMO Systems over Rayleigh-Product Channels

### 05 Jan 2024 v4

@@ C-D Proof of (19)
To establish the convergence rate of the cumulative density function of the MID, we first use Esseen inequality [53, p538] to conclude that there exists $C>0$ for any $T>0$ such that
$\sup\limits_{x\in\mathbb{R}}|\mathbb{P}(\sqrt{\frac{{Mn}}{V_{n}}}(I_{N,L,M}^{(n)}-\overline{C}(\sigma^{2}))\leq x)|\leq C\int_{0}^{T}u^{-1}|\Psi^{{\mathbb{W}},{\mathbb{Y}},{\mathbb{Z}}}_{norm}(u)-e^{-\frac{u^{2}}{2}}|\mathrm{d}u+\frac{C}{T}$ (151)
$\leq K(\int_{0}^{T}u^{-1}E(\frac{u}{\sqrt{V_{n}}},{\mathbb{A}})+\frac{1}{T}).$
Notice that the dominating of the error term in (147) is ${\mathcal{O}}(\frac{\mathcal{M}(1,u)+u}{\sqrt{M}})$ and for $T>1$,
@@
$\sup\limits_{x\in\mathbb{R}}|\mathbb{P}(\sqrt{\frac{{Mn}}{V_{n}}}(I_{N,L,M}^{(n)}-\overline{C}(\sigma^{2}))\leq x)|={\mathcal{O}}(n^{-\frac{1}{4}}).$ (153)
@@ Appendix D Proof of Theorem 3
$\mathbb{P}(\frac{\sqrt{Mn}(I_{N,L,M}^{(n+1)}-\overline{C}(\sigma^{2}))}{\sqrt{\hat{V}_{n+1}}}\leq z)\xrightarrow[]{n\xrightarrow[]{\rho,\eta,\kappa}\infty}\Phi(z),$ (154)
@@
$\mathbb{P}(\frac{\sqrt{Mn}(I_{N,L,M}^{(n+1)}-\overline{C}(\sigma^{2}))}{\sqrt{{V}_{n,+}}}\leq z)\xrightarrow[]{n\xrightarrow[]{\rho,\eta,\kappa}\infty}\Phi(z),$ (155)
where ${V}_{n,+}=V_{-}+\frac{\rho\Theta\operatorname{Tr}{\mathbb{A}}_{n+1}^{2}}{M}$. Therefore, we can obtain
$\mathbb{P}(\sqrt{Mn}(I_{N,L,M}^{(n+1)}-\overline{C}(\sigma^{2}))\leq r-\zeta)=\mathbb{P}(\frac{\sqrt{Mn}(I_{N,L,M}^{(n+1)}-\overline{C}(\sigma^{2}))}{\sqrt{{V}_{n,+}}}\leq\frac{r-\zeta}{\sqrt{{{V}_{n,+}}}})$ (156)
$\overset{(a)}{\geq}\begin{cases}\mathbb{P}(\frac{\sqrt{Mn}(I_{N,L,M}^{(n+1)}-\overline{C}(\sigma^{2}))}{\sqrt{V_{n,+}}}\leq\frac{r-\zeta}{\sqrt{V_{-}}})~{},r\leq 0\\ \mathbb{P}(\frac{\sqrt{Mn}(I_{N,L,M}^{(n+1)}-\overline{C}(\sigma^{2}))}{\sqrt{V_{n,+}}}\leq 0~{},r>0\end{cases}=\begin{cases}\Phi(\frac{r-\zeta}{\sqrt{V_{-}}})+c_{n},~{}~{}r\leq 0\\ 0.5+c_{n},~{}~{}r>0\end{cases}$
for the sequence $c_{n}\downarrow 0$. The inequality in step $(a)$ holds true since $V_{n,+}\geq V_{-}>0$ and the case for $r>0$ follows from the fact that $\frac{\operatorname{Tr}{\mathbb{A}}_{n+1}^{2}}{M}={\mathcal{O}}(M)$ such that $V_{n,+}\rightarrow\infty$. The lower bound can be obtained by taking the limit $\downarrow 0$.
@@
$\frac{\operatorname{Tr}{\mathbb{A}}^{2}_{n}}{M}=1-\frac{2\operatorname{Tr}{\mathbb{X}}^{(n)}({\mathbb{X}}^{(n)})^{H}}{Mn}+\frac{\operatorname{Tr}({\mathbb{X}}^{(n)}({\mathbb{X}}^{(n)})^{H})^{2}}{Mn^{2}}\xrightarrow[]{a.s.}\rho^{-1}.$ (158)
@@ D-A Proof of (19)
To establish the convergence rate of the upper and lower bounds, we will use the Esseen inequality in (151). The lower bound can be regarded as $\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M}=0$, which ${\mathbb{X}}$ can be constructed by orthogonal basis when $n>M$. In this case, the error term in (147) can be simplified as
$E(u,{\mathbb{A}})$ $={\mathcal{O}}(\frac{\mathcal{M}(V_{n},u)}{M}+\frac{V_{n}\mathcal{M}(V_{n},u)}{M^{2}}+\frac{u\mathcal{M}(V_{n},u)}{M}+\frac{u^{2}\mathcal{M}(V_{n},u)}{M^{2}}+\frac{u^{3}\mathcal{M}(V_{n},u)}{M^{3}}$ (159)
$+\frac{u}{M}+\frac{u\mathcal{M}(V_{n},u)}{M}+\frac{u^{2}\mathcal{M}(V_{n},u)}{M^{2}}+\frac{\mathcal{M}(V_{n},u)}{M}+(\frac{u}{M}+\frac{u^{2}}{M^{2}})\mathcal{M}(V_{n},u)).$
@@
$\sup\limits_{x\in\mathbb{R}}|\mathbb{P}(\sqrt{\frac{{Mn}}{V_{n}}}(I_{N,L,M}^{(n)}-\overline{C}(\sigma^{2}))\leq x)-\Phi(x)|={\mathcal{O}}(n^{-\frac{1}{2}}).$ (160)
@@
$\|\operatorname{Tr}{\mathbb{A}}^{4}\|\leq\|\operatorname{Tr}{\mathbb{I}}_{M}\|+4n^{-1}\|\operatorname{Tr}({\mathbb{X}}{\mathbb{X}}^{H})\|+6n^{-2}\|\operatorname{Tr}({\mathbb{X}}{\mathbb{X}}^{H})^{2}\|+4n^{-3}\|\operatorname{Tr}({\mathbb{X}}{\mathbb{X}}^{H})^{3}\|+n^{-4}\|\operatorname{Tr}({\mathbb{X}}{\mathbb{X}}^{H})^{4}\|={\mathcal{O}}(M).$ (161)
@@
$E(u,{\mathbb{A}})$ $={\mathcal{O}}(\frac{\mathcal{M}(V_{n},u)}{M}+\frac{V_{n}\mathcal{M}(V_{n},u)}{M^{2}}+\frac{u\mathcal{M}(V_{n},u)}{M}+\frac{u^{2}\mathcal{M}(V_{n},u)}{M^{2}}+\frac{u^{3}\mathcal{M}(V_{n},u)}{M^{3}}$ (162)
$+\frac{u}{M}+\frac{u\mathcal{M}(V_{n},u)}{M}+\frac{u^{2}\mathcal{M}(V_{n},u)}{M^{2}}\frac{\mathcal{M}(V_{n},u)}{M}+[\frac{1}{M}+\frac{u}{M}+\frac{u^{2}}{M^{2}}]\mathcal{M}(V_{n},u)),$
@@
$\sup\limits_{x\in\mathbb{R}}|\mathbb{P}(\sqrt{\frac{{Mn}}{V_{n}}}(I_{N,L,M}^{(n)}-\overline{C}(\sigma^{2}))\leq x)-\Phi(x)|={\mathcal{O}}(n^{-\frac{1}{2}}).$ (163)
@@ Appendix F Proof of Proposition 2
$\overline{\omega}^{4}(\frac{\kappa\omega^{2}(1+\delta\overline{\omega})}{1+\delta\overline{\omega}^{2}}-\frac{\kappa\omega\omega^{\prime}}{\delta(1+\delta\overline{\omega}^{2})})$ (167)

### 09 Apr 2024 v5

@@ C-D Proof of (19)
To establish the convergence rate for the CDF of the MID, we will use Esseen inequality [53, p538], which says that there exists $C>0$ for any $T>0$ such that
$\sup\limits_{x\in\mathbb{R}}|\mathbb{P}\left(\sqrt{\frac{{Mn}}{V_{n}}}(I_{N,L,M}^{(n)}-\overline{C}(\sigma^{2})\right)\leq x)|\leq C\int_{0}^{T}u^{-1}|\Psi^{{\mathbb{W}},{\mathbb{Y}},{\mathbb{Z}}}_{norm}(u)-e^{-\frac{u^{2}}{2}}|\mathrm{d}u+\frac{C}{T}$ (151)
$\leq K(\int_{0}^{T}u^{-1}E\left(\frac{u}{\sqrt{V_{n}}},{\mathbb{A}})+\frac{1}{T}\right).$
Notice that the dominating term in (147) is ${\mathcal{O}}(\frac{\mathcal{M}(u)+u}{\sqrt{M}})$ and for $T>1$,
@@
$\sup\limits_{x\in\mathbb{R}}|\mathbb{P}\left(\sqrt{\frac{{Mn}}{V_{n}}}(I_{N,L,M}^{(n)}-\overline{C}(\sigma^{2}))\leq x\right)|={\mathcal{O}}(n^{-\frac{1}{4}}),$ (153)
@@ Appendix D Proof of Theorem 3
$\mathbb{P}\Biggl{(}\frac{\sqrt{Mn}(I_{N,L,M}^{(n+1)}-\overline{C}(\sigma^{2}))}{\sqrt{\hat{V}_{n+1}}}\leq z\Biggr{)}\xrightarrow[]{n\xrightarrow[]{\rho,\eta,\kappa}\infty}\Phi(z),$ (154)
@@
$\mathbb{P}\Biggl{(}\frac{\sqrt{Mn}(I_{N,L,M}^{(n+1)}-\overline{C}(\sigma^{2}))}{\sqrt{{V}_{n,+}}}\leq z\Biggr{)}\xrightarrow[]{n\xrightarrow[]{\rho,\eta,\kappa}\infty}\Phi(z),$ (155)
where ${V}_{n,+}=V_{-}+\frac{\rho\Theta\operatorname{Tr}({\mathbb{A}}_{n+1}^{2})}{M}$. Therefore, we can obtain
$\mathbb{P}(\sqrt{Mn}(I_{N,L,M}^{(n+1)}-\overline{C}(\sigma^{2}))\leq r-\zeta)=\mathbb{P}\Biggl{(}\frac{\sqrt{Mn}(I_{N,L,M}^{(n+1)}-\overline{C}(\sigma^{2}))}{\sqrt{{V}_{n,+}}}\leq\frac{r-\zeta}{\sqrt{{{V}_{n,+}}}}\Biggr{)}$ (156)
$\overset{(a)}{\geq}\begin{cases}\mathbb{P}\Bigl{(}\frac{\sqrt{Mn}(I_{N,L,M}^{(n+1)}-\overline{C}(\sigma^{2}))}{\sqrt{V_{n,+}}}\leq\frac{r-\zeta}{\sqrt{V_{-}}}\Bigr{)}~{},r\leq 0\\ \mathbb{P}\Bigl{(}\frac{\sqrt{Mn}(I_{N,L,M}^{(n+1)}-\overline{C}(\sigma^{2}))}{\sqrt{V_{n,+}}}\Bigr{)}\leq 0~{},r>0\end{cases}=\begin{cases}\Phi(\frac{r-\zeta}{\sqrt{V_{-}}})+c_{n},~{}~{}r\leq 0\\ 0.5+c_{n},~{}~{}r>0\end{cases}$
for the sequence $c_{n}\downarrow 0$. The inequality in step $(a)$ holds true since $V_{n,+}\geq V_{-}>0$ and the case for $r>0$ follows from the fact that $\frac{\operatorname{Tr}({\mathbb{A}}_{n+1}^{2})}{M}={\mathcal{O}}(M)$ such that $V_{n,+}\rightarrow\infty$. The lower bound can be obtained by taking the limit $\downarrow 0$.
@@
$\frac{\operatorname{Tr}({\mathbb{A}}^{2}_{n})}{M}=1-\frac{2\operatorname{Tr}({\mathbb{X}}^{(n)}({\mathbb{X}}^{(n)})^{H})}{Mn}+\frac{\operatorname{Tr}(({\mathbb{X}}^{(n)}({\mathbb{X}}^{(n)})^{H})^{2})}{Mn^{2}}\xrightarrow[]{a.s.}\rho^{-1}.$ (158)
@@ D-A Proof of (19)
To establish the convergence rate of the upper and lower bounds, we will use the Esseen inequality in (151). The lower bound can be regarded as $\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M}=0$, which ${\mathbb{X}}$ can be constructed by orthogonal basis when $n>M$. In this case, the error term in (147) can be simplified as
$E(u,{\mathbb{A}})$ $={\mathcal{O}}\Bigl{(}\frac{\mathcal{M}(V_{n},u)}{M}+\frac{V_{n}\mathcal{M}(V_{n},u)}{M^{2}}+\frac{u\mathcal{M}(V_{n},u)}{M}+\frac{u^{2}\mathcal{M}(V_{n},u)}{M^{2}}+\frac{u^{3}\mathcal{M}(V_{n},u)}{M^{3}}$ (159)
$+\frac{u}{M}+\frac{u\mathcal{M}(V_{n},u)}{M}+\frac{u^{2}\mathcal{M}(V_{n},u)}{M^{2}}+\frac{\mathcal{M}(V_{n},u)}{M}+(\frac{u}{M}+\frac{u^{2}}{M^{2}})\mathcal{M}(V_{n},u)\Bigr{)}.$
@@
$\sup\limits_{x\in\mathbb{R}}\Bigl{|}\mathbb{P}\Bigl{(}\sqrt{\frac{{Mn}}{V_{n}}}(I_{N,L,M}^{(n)}-\overline{C}(\sigma^{2}))\leq x\Bigr{)}-\Phi(x)\Bigr{|}={\mathcal{O}}(n^{-\frac{1}{2}}).$ (160)
@@
$\|\operatorname{Tr}({\mathbb{A}}^{4})\|\leq\|\operatorname{Tr}({\mathbb{I}}_{M})\|+4n^{-1}\|\operatorname{Tr}({\mathbb{X}}{\mathbb{X}}^{H})\|+6n^{-2}\|\operatorname{Tr}(({\mathbb{X}}{\mathbb{X}}^{H})^{2})\|+4n^{-3}\|\operatorname{Tr}(({\mathbb{X}}{\mathbb{X}}^{H})^{3})\|+n^{-4}\|\operatorname{Tr}(({\mathbb{X}}{\mathbb{X}}^{H})^{4})\|={\mathcal{O}}(M).$ (161)
@@
$E(u,{\mathbb{A}})$ $={\mathcal{O}}\Bigl{(}\frac{\mathcal{M}(V_{n},u)}{M}+\frac{V_{n}\mathcal{M}(V_{n},u)}{M^{2}}+\frac{u\mathcal{M}(V_{n},u)}{M}+\frac{u^{2}\mathcal{M}(V_{n},u)}{M^{2}}+\frac{u^{3}\mathcal{M}(V_{n},u)}{M^{3}}$ (162)
$+\frac{u}{M}+\frac{u\mathcal{M}(V_{n},u)}{M}+\frac{u^{2}\mathcal{M}(V_{n},u)}{M^{2}}\frac{\mathcal{M}(V_{n},u)}{M}+[\frac{1}{M}+\frac{u}{M}+\frac{u^{2}}{M^{2}}]\mathcal{M}(V_{n},u)\Bigr{)},$
@@
$\sup\limits_{x\in\mathbb{R}}\Bigl{|}\mathbb{P}\Bigl{(}\sqrt{\frac{{Mn}}{V_{n}}}(I_{N,L,M}^{(n)}-\overline{C}(\sigma^{2}))\leq x\Bigr{)}-\Phi(x)\Bigr{|}={\mathcal{O}}(n^{-\frac{1}{2}}).$ (163)
@@ Appendix F Proof of Proposition 2
$\overline{\omega}^{4}\Bigl{(}\frac{\kappa\omega^{2}(1+\delta\overline{\omega})}{1+\delta\overline{\omega}^{2}}-\frac{\kappa\omega\omega^{\prime}}{\delta(1+\delta\overline{\omega}^{2})}\Bigr{)}$ (167)